Find the center and radius of the circle with the given equation.
x^2+y^2-8y=9

Found 2 solutions by fractalier, Cromlix:
Answer by fractalier(6550)

We have to complete the square by adding 16 to both sides...from
x^2+y^2-8y=9
x^2+y^2-8y+16 = 25
x^2 + (y-4)^2 = 5^2
so that the center is at
(0,4) with a radius of 5.

Answer by Cromlix(4381)

Hi there,
x^2+y^2-8y=9
x^2 + y^2 - 8y - 9 = 0
Centre is obtained by halving the integers before
the x and y components and making them negative
of the equation
Centre (0, 4)
...........
Radius = sqrt x coord of centre squared +
y coord of centre squared - the integer at
the end of the equation (-9)
Radius = sqrt of 0^2 + 4^2 + 9
= sqrt25
= 5
Hope this helps :-)